# Have Faith in Faithfulness: Going Beyond Circuit Overlap When Finding Model Mechanisms

### v1

@@ 6 Discussion
In this study, we have analyzed the question of faithfulness in circuit-finding, discovering that while EAP promised to automatically find circuits, its faithfulness was poor compared to EAP-IG. However, faithfulness is not the only desideratum of circuits: while measuring faithfulness ensures that our circuits stray little from whole-model performance, it does not guarantee that they are complete, i.e. not missing any important (negative-acting) components. We aim to avoid this problem by selecting components based on absolute magnitude, thus including both positively and negatively important components. However, future work could study circuits through the lens of completeness, as in Wang et al. (2023).
We also found that while zero and full overlap predict low and high cross-task faithfulness, moderate overlap predicts little. This occurs largely because cross-task faithfulness depends not on overlap generally, but on which specific nodes or edges overlap (e.g. the edge from the input to MLP 0 is often crucial). This issue might be avoided with a metric, such as weighted graph edit distance, that takes into account each edge’s importance (score). This would require published circuits to report edge scores, a practice that we would encourage.
To conclude, as circuits become more popular and easier to find, practitioners’ research questions increasingly involve not single circuits but circuits across tasks and models. This is an important development—but as we attempt to answer such questions, we must be careful not to be misled about the actual mechanisms that underlie model behavior across tasks; these mechanisms are the core of circuits research. Measuring cross-task faithfulness and characterizing component behavior across tasks, will yield sounder results.
@@ Acknowledgments
This research has been supported by an AI Alignment grant from Open Philanthropy, the Israel Science Foundation (grant No. 448/20), and an Azrieli Foundation Early Career Faculty Fellowship.
@@ Appendix A Dataset and Metric Details
Here, we provide additional details regarding our datasets. The Hypernymy task formulation was inspired by Hearst’s (1992) work on automatically acquiring hypernyms from corpora. By using prompts that resemble contexts in which hypernyms naturally appear in data, we are able to elicit hypernym-retrieval behavior from even weaker language models. Similar strategies have been used by later work on hypernymy in masked language models (Ettinger, 2020; Ravichander et al., 2020; Hanna & Mareček, 2021). The Country-Capital task is also one that has been studied previously, though it has generally not the focus of any one work (see Appendix H of Merullo et al. (2024)).
@@
Country-Capital logit diff 6.12 -6.72
@@ Appendix C EAP-IG Requires Few Steps
Appendix D Circuit Search
@@
Appendix E Approximation Error and Kendall Correlations
@@
Appendix F Within-Task Overlap and Faithfulness
@@
Appendix G Circuit Overlap Significance Testing
@@
Appendix H Greater-Than Variants
Refer to captionRefer to caption
Figure 10: Faithfulness plots for Greater-Than (Price) and Greater-Than (Sequence).
In this section, we discuss the two Greater-Than variants, Greater-Than (Price) and Greater-Than (Sequence). Before running our overlap and cross-task faithfulness experiments on them, we confirm that faithful circuits for them do indeed exist; in Figure 10, we see that this is the case. For the overlap and cross-task faithfulness experiments, we take the 300 and 200 edge EAP-IG-KL circuits for Greater-Than (Price) and Greater-Than (Sequence) respectively.
Appendix I Asymmetric Overlap Measures
@@
Appendix J Replications in Larger Models

### v2

@@ 6 Discussion
In this study, we have analyzed the question of faithfulness in circuit-finding, discovering that while EAP promised to automatically find circuits, its circuits’ faithfulness was poor compared to EAP-IG’s. We also found that while zero and full overlap predict low and high cross-task faithfulness, moderate overlap predicts little, leading us to recommend testing faithfulness separately from overlap. However, many open questions remain; here, we discuss these open questions and recommend best practices for circuit finding.
Completeness Faithfulness is not the only potential circuit desideratum: while measuring faithfulness ensures that our circuits stray little from whole-model performance, it does not guarantee that they are complete, i.e. not missing any important components, even those that inhibit model performance. Notably, while many circuits may achieve a given faithfulness, a circuit that is both faithful and complete should be relatively unique. We aimed to find complete circuits by selecting edges based on absolute score magnitude, thus including both positively and negatively important edges; this contrasts with other work, which prioritized finding positively important components (Prakash et al., 2024).
We would like to prove that EAP-IG’s circuits are complete, but testing completeness is challenging. Wang et al. (2023) compare compare model and circuit behavior under random ablations, reasoning that a complete circuit should remain similar to the whole model even under arbitrary ablations. However, this is costly, and it is unclear how to choose sensible random ablations. If manually-found circuits are taken as ground-truth complete circuits, then circuit overlap could be considered metric of completeness, rather than faithfulness. Still, overlap as a metric is flawed: a completeness metric should ideally, like faithfulness, be measurable without a ground truth, as manually-found “ground-truth” circuits may be unreliable.
Briding the gap between overlap and faithfulness Overlap and cross-task faithfulness yield different results. This occurs in large part because the latter depends not on overlap generally, but on which specific nodes or edges overlap: for example, the edge from the input to MLP 0 is often crucial. This issue might be avoided with a metric, such as weighted graph edit distance, that takes into account each edge’s importance (score). This would require published circuits to report edge scores, a practice that we would encourage.
Best practices for circuits As circuits become more popular and easier to find, practitioners’ research questions increasingly involve comparing circuits across tasks and models. This is an important development, but as we attempt to answer such questions, we must be careful not to be misled about the actual mechanisms that underlie model behavior across tasks; these mechanisms are the core of circuits research. Measuring cross-task faithfulness and characterizing component behavior across tasks, will yield sounder results.
@@ Acknowledgments
The authors thank Neel Nanda for suggesting the Clean-Corrupted baseline and other comparisons. MH is partially supported by an OpenAI Superalignment Fellowship. This research has been supported by an AI Alignment grant from Open Philanthropy, the Israel Science Foundation (grant No. 448/20), and an Azrieli Foundation Early Career Faculty Fellowship.
@@ Appendix A Dataset and Metric Details
Here, we provide additional details regarding our datasets. The Hypernymy task formulation was inspired by Hearst’s (1992) work on automatically acquiring hypernyms from corpora. By using prompts that resemble contexts in which hypernyms naturally appear in data, we are able to elicit hypernym-retrieval behavior from even weaker language models. Similar strategies have been used by later work on hypernymy in masked language models (Ettinger, 2020; Ravichander et al., 2020; Hanna & Mareček, 2021). The Capital-Country task is also one that has been studied previously, though it has generally not the focus of any one work (see Appendix H of Merullo et al. (2024)).
@@
Capital-Country logit diff 6.12 -6.72
@@ Appendix C EAP-IG Requires Few Steps
Appendix D Comparing EAP-IG To Other Variants
Although we introduce one particular way of implementing EAP-IG in this paper, contemporary work has proposed many alternatives. In this section, we discuss a number of variants on EAP-IG, and then compare their performance on the tasks studied throughout this paper. We find that our version of EAP-IG outperforms other EAP-IG variants, though one baseline in particular is quite strong.
D.1 EAP-IG Variants and Baselines
EAP-IG in activation space (Marks et al., 2024)
Marks et al. search for circuits of features, rather than nodes and edges. As a result, their conception of the transformer’s graph is somewhat different from ours; they break the input embeddings, each block of attention heads, each MLP, and each point in the residual stream after an MLP, into a thousands of sparse features, each of which constitutes a node. Unlike our work, feature nodes are not connected to all downstream features nodes; instead, they are connected only to nodes in the next layer.
Marks et al. propose an EAP-IG method to score the edges of the thousands of feature nodes efficiently. However, to find the score of an edge $(u,v)$, they do not average the gradient when interpolating between different inputs with embeddings $z$ and $z^{\prime}$, Rather, they take the gradient of $v$ when setting the activations of $u$ to a combination of its clean $u$ and corrupted outputs. Formally, they compute the score of an edge as:
$(z^{\prime}_{u}-z_{u})\frac{1}{m}\sum_{k=1}^{m}\frac{\partial L(s|\text{do}(z_{u,out}=z_{u,corrupted}+\frac{k}{m}(z_{u,clean}-z_{u,corrupted}))}{\partial z_{v}}.$ (4)
This quantity cannot be calculated with just $m$ forward passes; performing the $do$ operation on multiple nodes that are not calculated in parallel will yield incorrect scores. As a result, Marks et al. iterate over sublayers that are calculated in parallel (the aforementioned input embeddings, attention head blocks, MLPs, and residual stream points), and perform $m$ forward passes for each. In the context of feature circuits, this is a significant improvement over ablating features individually; however, in the context of component / edge circuits, this scales worse than EAP-IG in input space.
Partial EAP-IG in activation space (Miller et al., 2024)
AutoCircuit, the circuit-finding framework introduced by Miller et al. (2024), introduces its own EAP-IG variant, called mask_gradient_prune_scores. Though Miller et al. do not discuss it directly, an analysis of their code suggests that this is indeed a distinct EAP-IG variant. They compute the score of an edge ($u,v$) as roughly:
$(z^{\prime}_{u}-z_{u})\frac{1}{m}\sum_{k=1}^{m}\frac{\partial L(s|\text{do}(z_{u}=z_{u}^{\prime}+\frac{k}{m}(z_{u}-z_{u}^{\prime}))}{\partial z_{v}}.$ (5)
In contrast to the prior formulation, which set the output of each node to the interpolation between its clean and corrupted outputs, this formulation sets each node’s output to an interpolation of its current and corrupted outputs. To calculate this quantity exactly, the $do$ operation should only be performed in parallel on independent nodes, much like in the prior formulation; however, Miller et al. compute this in parallel for all nodes. Though the implementation we test is identical to theirs, our mathematical formulation of their procedure is thus imprecise. Letting $V$ be the set of all non-logit nodes in the model’s computational graph, we could instead formulate it as follows:
$(z^{\prime}_{u}-z_{u})\frac{1}{m}\sum_{k=1}^{m}\frac{\partial L(s|\text{do}(\forall n\in V:z_{n}=z_{n}^{\prime}+\frac{k}{m}(z_{n}-z_{n}^{\prime}))}{\partial z_{v}}.$ (6)
Clean-Corrupted Baseline
In our definition of EAP-IG, we average gradients over an interpolation between clean and corrupted inputs. However, Appendix C suggests that the interpolated steps are not very important, as EAP-IG works with $m$ as low as 3. A reasonable baseline is thus EAP-IG, where the gradients used in scoring are just the average of the gradients on clean and corrupted inputs; this is distinct from EAP-IG when $m=2$, as EAP-IG never calculates gradients using fully corrupted inputs.
$(z^{\prime}_{u}-z_{u})^{\top}\left(\frac{1}{2}\nabla_{v}L(s)+\frac{1}{2}\nabla_{v}L(s^{\prime})\right)$ (7)
D.2 Results
Figure 8: Faithfulness of circuits found using EAP-IG and variants.
We run EAP, EAP-IG (with 2 and 5 steps), activation patching and all three methods mentioned above on all 6 tasks that we study, as in Section 4. Our results (Figure 8) show that our EAP-IG works roughly as well as Marks et al.’s slower EAP-IG variant across tasks. It also outperforms Miller et al.’s EAP-IG variant on all tasks but IOI. However, the Clean-Corrupted baseline is quite strong: EAP-IG outperforms it only on the SVA task. This is not especially surprising, given how few steps EAP-IG needed to succeed, but EAP-IG with $m=2$ steps does underperform EAP-IG with 5 steps and the Clean-Corrupted baseline. While our EAP-IG has a slight edge on SVA, the Clean-Corrupted baseline seems like a strong method in its own right, and certainly outperforms EAP.
Appendix E Circuit Search
@@
Appendix F Approximation Error and Kendall Correlations
@@
Appendix G Within-Task Overlap and Faithfulness
@@
Appendix H Circuit Overlap Significance Testing
@@
Appendix I Greater-Than Variants
Figure 11: Faithfulness plots for Greater-Than (Price) Greater-Than (Sequence), and Country-Capital.
In this section, we discuss the two Greater-Than variants, Greater-Than (Price) and Greater-Than (Sequence), and the Country-Capital task. Before running our overlap and cross-task faithfulness experiments on them, we confirm that faithful circuits for them do indeed exist; in Figure 11, we see that this is the case. For the overlap and cross-task faithfulness experiments, we take the 300, 200, and 300 edge EAP-IG-KL circuits for Greater-Than (Price) and Greater-Than (Sequence), and Country-Capital respectively.
Appendix J Asymmetric Overlap Measures
@@ Appendix I Greater-Than Variants
Appendix K Replications in Larger Models